# MIPHEI-vit: Multiplex Immunofluorescence Prediction from H&E Images using ViT Foundation Models

### v1

@@ 6 Results
This section presents an ablation study to identify critical model components, comparisons with the state-of-the-art HEMIT method, and a detailed analysis of marker-specific performance. We selected HEMIT for benchmarking as it addresses a similar problem, provides public code and data for direct comparison, and outperforms baselines like Pix2Pix used in other related work such as SHIFT.
Figure 5: Evaluation analysis: (a) Orion test set: Performance comparison across all Orion markers between MIPHEI, HEMIT*, HEMIT (all trained on Orion train set) and a stratified random model predicting classes based on cell type proportions. Markers are grouped by general functions, with hierarchical relationships indicated by arrows. Cell classification model is a logistic regression model trained on Orion validation cells. (b) HEMIT dataset: Comparison between MIPHEI (trained on Orion), the HEMIT model (trained on HEMIT train set) and a stratified random model. Cell classification model is a logistic regression: trained on 5% of training cells for MIPHEI, and all available training cells for the HEMIT model. (c) IMMUcan (consecutive sections): Pearson correlation and regression plots (with linear fits) between predicted and pseudo-labeled cell type counts from 17k tiles using MIPHEI.
@@ 6.1 Ablation studies
Table 1: Ablation Study: Performance Metrics Across Model Configurations on ORION test set
Configuration PSNR SSIM Cell AUC Cell F1
Impact of GAN Discriminator (UNETR H-optimus-0 LoRA)
Generator only 27.86 0.840 0.868 0.431
Pix2Pix (GAN) 27.00 0.830 0.817 0.410
Impact of Foundation Model Encoder (UNETR with LoRA)
CTransPath 26.96 0.837 0.812 0.351
Univ2 27.73 0.838 0.862 0.424
H-optimus-0 (selected) 27.86 0.840 0.868 0.431
Encoder Finetuning Method (UNETR H-optimus-0)
Frozen encoder 27.44 0.836 0.857 0.413
LoRA (selected) 27.86 0.840 0.868 0.431
@@
U-Net-ResNet50 26.54 0.832 0.812 0.344
U-Net-ConvNeXtv2 Large 27.40 0.839 0.840 0.379
HEMIT* 27.08 0.837 0.831 0.360
UNETR 27.86 0.840 0.868 0.431
MIPHEI (Best) 27.78 0.837 0.876 0.438
@@ 6.1.1 Impact of Discriminator
6.1.2 Model Architecture
We next compared four U-Net-inspired architectures for our mIF prediction model: two convolution-based encoders pre-trained on ImageNet—ResNet50 (37M parameters) and ConvNeXt v2 Large (203M parameters)—and two U-Net variants that both use the H-optimus-0 foundation model as encoder but differ in their ViT integration strategies—UNETR (1.17B parameters, 36M trainable via LoRA) and ViTMatte (1.14B parameters, 6.6M trainable via LoRA). U-Net with a ResNet50 encoder showed lower performance, likely due to its smaller size and lack of domain-specific pretraining. In contrast, U-Net-ConvNeXt achieved strong results, benefiting from greater model capacity and attention-like mechanisms. Transformer-based variants using H-optimus-0 encoder significantly outperformed convolutional ones, emphasizing the advantage of employing pretrained foundation models for the encoder component in our task. UNETR and ViTMatte achieved similar performance, but ViTMatte converged much faster, reaching UNETR’s 10-epoch performance after the first one only. This is likely due to its convolutional ’detail capture’ stream, which provides intermediate skip features that complement ViT representations and aid spatial reconstruction more effectively than UNETR’s token-based upsampling. Overall, ViTMatte offers the best trade-off between performance and training efficiency.
@@ 6.1.4 Finetuning Strategies
6.2 Comparison with State of the Art (HEMIT model)
To benchmark our proposed model, MIPHEI, against the state-of-the-art HEMIT, we evaluate cross-dataset generalization using five models: (1) MIPHEI, trained on ORION, (2) the official HEMIT checkpoint trained on the HEMIT dataset, (3) HEMIT* trained on ORION using the HEMIT-architecture, but with the MIPHEI training scheme (weighted MSE loss, no discriminator), (4) HEMIT-Orion, trained on ORION using the original HEMIT pipeline, and (5) a stratified random model as a baseline, assigning cell type probabilities based on their distribution in the test datasets, with metrics averaged over 100 runs to account for variability in random sampling. Since most cell subtypes are not identifiable from H&E alone by a pathologist, the task is challenging, and the random baseline provides a reference to show that our model outperforms chance-level expectations. Evaluation is conducted on the ORION test set (in-domain for models trained on ORION), on the HEMIT validation and test sets and the IMMUcan dataset, both serving as external out-of-domain benchmarks.
6.2.1 Evaluation on ORION test set
Table 2: Overall Cell-Level Performances on Test Sets
Cell Metric MIPHEI (Orion) HEMIT (HEMIT) HEMIT* (Orion) HEMIT (Orion) Random
ORION Dataset (Test Set)
AUC 0.876 - 0.831 0.701 -
F1 0.438 - 0.360 0.253 0.140
HEMIT Dataset (Average Validation & Test Sets)
AUC 0.844 0.863 0.764 0.598 -
F1 0.701 0.663 0.471 0.481 0.333
IMMUcan Dataset
Pearson 0.690 - 0.667 0.422 -
The results of performance evaluation of all tested models is shown in Table 2, Figure 5). HEMIT-Orion, trained with the original HEMIT pipeline, outperformed the random model, but showed the lowest performance across all markers except PD-L1, for which all methods performed poorly. We hypothesize that this was due to the use of the L1 loss, which may be less effective for markers with low intensity or high background proportion, as suggested by the stronger results of HEMIT* trained with our setup. On the other hand, MIPHEI achieved the best overall performance, surpassing both HEMIT* and HEMIT-Orion across nearly all markers. The only exception was Ki67, where HEMIT* performed slightly better, achieving a Cell F1 score 0.3% higher than MIPHEI.
These results confirm that MIPHEI provides the most robust predictions within its training domain. The improved performance of HEMIT* over HEMIT-Orion further demonstrates the effectiveness of our training pipeline, even when applied to an alternative architecture.
6.2.2 Evaluation on HEMIT dataset
The HEMIT dataset serves as an external test set for assessing the robustness of our approach. We compare MIPHEI (train: ORION), with the HEMIT model (train: HEMIT, released checkpoint). As Figure 3.B suggests, the HEMIT validation and test sets come from different domains, but both are contained within the training domain, with the test domain being more strongly represented. Hence, we report cell-level performances separately for each. We use logistic regression as the cell classifier, trained on 5% of the training cells for MIPHEI and all training cells for HEMIT. As MIPHEI predicts 15 ORION markers while HEMIT predicts only Pan-CK and CD3, each model’s classifier is trained on the full set of markers predicted by its respective model.
On the HEMIT test set, MIPHEI slightly outperforms HEMIT by +3% F1 score for Pan-CK (Figure 5), but HEMIT significantly outperforms MIPHEI on CD3, with a +22% Cell F1 score. However, on the HEMIT validation set, MIPHEI surpasses HEMIT for both markers, achieving +18% F1 for Pan-CK and +17% for CD3. This is a strong result, as MIPHEI was evaluated on an external dataset with only minimal adaptation of the auxiliary cell classification stage, using just 5% of the training data. In contrast, the HEMIT model was trained on data that included images from both the validation and test domains of the HEMIT dataset (Figure 3.C). MIPHEI demonstrates stronger generalization, likely due to its foundation model encoder, enabling better adaptation to unseen domains.
6.2.3 Evaluation on IMMUcan dataset
We further evaluate model generalization on the IMMUcan dataset by performing a correlation analysis between predicted cell type counts from H&E and pseudo-label counts from consecutive mIF sections. Pearson correlation is computed over 0.26 mm² regions for MIPHEI, HEMIT*, and HEMIT. Each model uses its own logistic regression cell classifier, trained on ORION validation cells for MIPHEI and HEMIT*, and on all HEMIT training cells for HEMIT.
Across all markers, MIPHEI achieves the highest Pearson correlation, outperforming both HEMIT* and HEMIT, further confirming its superiority in cross-domain generalization (Figure 5).
@@ 6.3.1 In-domain performance analysis
Markers with the highest performance include epithelial markers E-cadherin (F1 0.903) and Pan-CK (F1 0.884), which label epithelial cells forming well-defined clusters and glandular structures in H&E. The immune marker CD45 (F1 0.681) also performs well. While it mainly identifies lymphocytes, which are easily recognizable in H&E, it also includes cells like monocytes, which are harder to spot, making prediction a bit more challenging.
Markers with moderate performance included SMA (F1 0.564), which labels smooth muscle cells, and CD31 (F1 0.386), which marks endothelial cells. Their more subtle morphological features may have contributed to lower accuracy: smooth muscle cells have a spindle-shaped appearance but can be confused with fibroblasts, while endothelial cells are typically sparse and located within vessel structures, making them harder to distinguish. Immune subtype markers CD3e, CD45RO, CD4, CD8a, and CD20 (F1 0.229–0.572) showed moderate performance, with broader T-cell markers like CD3e achieving better performance than more specific ones like CD8a, which are harder to identify. While lymphocytes are visible in H&E, their subtypes remain indistinguishable to pathologists, highlighting our model’s value on this difficult task. Macrophage markers CD68 (F1 0.362) and CD163 (F1 0.206) faced similar challenges, as their heterogeneity complicates identification, with CD163 being a macrophage subtype of CD68, further complicating prediction. Markers with the lowest performance are FOXP3 (F1 0.114) and PD-L1 (F1 0.048), both challenging for different reasons. FOXP3 marks rare regulatory T-cells, a highly specific CD4+ subtype, making it one of the most difficult immune markers to predict. PD-L1, a functional marker, lacks a clear association with a single cell type and shows irregular expression across various cells, making probably prediction from H&E nearly impossible.
In summary, structural markers and broad immune markers seem predictable with high accuracy. Markers defining specific immune subtypes are harder to predict, and some functional markers are not predictable with an accuracy high enough to be applicable in practice.
@@ 6.3.2 Out-of-domain performance analysis
Our external validation on HEMIT and IMMUcan confirms that MIPHEI, trained on ORION, generalizes well to datasets with domain shifts in mIF technology and H&E appearance, requiring minimal adaptation.
On the HEMIT dataset, analysis is limited by the small number of available markers, but our model performs well with only minimal adaptation of the cell classifier using 5% of the training cells.
On the IMMUcan dataset, we observe strong correlation for CD4 (0.80) CD3e (0.73) and CD8 (0.69), and moderate correlation for FOXP3 (0.60) and Pan-CK (0.63). We also observe overdetection for CD3e (i.e. cells with CD3e prediction and measured 0 expression) and underdetection of Pan-CK (cells with predicted 0-expression and positive according to the measurement). Manual inspection showed that this was mainly due to low-quality tiles with high auto-fluorescence.

### v2

@@ 6 Results
This section presents an ablation study to identify critical model components, a comparison with other methods, including HEMIT, ROSIE, DiffusionFT, and Pix2Pix, and a detailed analysis of marker-specific performance. We exclude HistoPlexer (Andani et al., 2025), as it is designed for consecutive-section supervision, which is not relevant in our same-section evaluation setting.
@@ 6.1 Ablation studies
Configuration Reconstruction Cell Classification
PSNR SSIM Pearson AUPRC F1
GAN Discriminator (MIPHEI-UNETR H-optimus-0 LoRA)
Pix2Pix (GAN) 30.88 0.942 0.395 0.531 0.499
Generator only 31.89 0.951 0.466 0.564 0.515
Foundation Model Encoder (MIPHEI-UNETR with LoRA)
CTransPath 30.77 0.947 0.357 0.472 0.449
UNIv2 31.88 0.950 0.454 0.554 0.508
H-optimus-0 (selected) 31.89 0.951 0.466 0.564 0.515
Encoder Finetuning Method (MIPHEI-UNETR H-optimus-0)
Frozen encoder 31.63 0.949 0.434 0.541 0.497
LoRA (selected) 31.89 0.951 0.466 0.564 0.515
@@
MIPHEI-ConvNeXtv2 L 31.51 0.949 0.404 0.508 0.473
MIPHEI-HEMIT 31.38 0.949 0.388 0.492 0.457
MIPHEI-UNETR 31.89 0.951 0.466 0.564 0.515
MIPHEI (Best) 31.93 0.951 0.470 0.574 0.523
Table 2: Ablation study on the OrionCRC test set with the optimal MIPHEI configuration shown in bold. Reconstruction performance is evaluated at the pixel level using PSNR, SSIM, and Pearson correlation. Cell-level performance is also reported using AUPRC and F1 score, as described in Section 5.2.2.
@@ 6.1.1 Impact of Discriminator
6.1.2 Architecture Comparison
We compared three MIPHEI variants for mIF prediction. One uses convolutional encoder pretrained on ImageNet: ConvNeXt v2 Large (203M parameters). The other two variants use the H-optimus-0 foundation model as encoder but differ in their ViT integration strategies: UNETR (1.17B parameters, 36M trainable via LoRA) and ViTMatte (1.14B parameters, 6.6M trainable via LoRA). MIPHEI-ConvNeXt achieved strong results, benefiting from high model capacity and attention-like mechanisms. Transformer-based variants using H-optimus-0 encoder significantly outperformed the convolutional one, emphasizing the advantage of employing pretrained foundation models for the encoder component in our task. UNETR and ViTMatte achieved similar performance, but ViTMatte converged much faster, reaching UNETR’s 10-epoch performance after the first one only. This is likely due to its lightweight convolutional Detail Capture module, which more directly complements ViT representations and aid spatial reconstruction compared to UNETR’s token-based upsampling. Overall, ViTMatte offers the best trade-off between performance and training efficiency and was therefore selected as the default MIPHEI architecture.
@@ 6.1.4 Finetuning Strategies
6.2 Benchmark with Existing Methods
Figure 6: Architectures used in our benchmark. Inference pipelines of the compared models, including CNN-based (top left), diffusion-based (top right), and U-Net-inspired architectures (bottom). Blue denotes convolutional blocks, yellow input images, purple predictions, and green intermediate tensors.
To benchmark MIPHEI, against existing approaches, we evaluate in-domain and cross-dataset generalization using the following models: (1) MIPHEI; (2) MIPHEI-ConvNeXt, a variant using a ConvNeXt encoder; (3) HEMIT; (4) ROSIE; (5) DiffusionFT; (6) a Nuclear Morphometry simple baseline relying on nucleus shape and H&E-derived statistics (e.g., area, orientation, hematoxylin intensity); (7) an Upper Bound model using real mIF-derived single-cell expressions, restricted to the intersection of markers available in each dataset and the predicted marker panel;(8) and a prevalence-base random model, which assigns cell-type probabilities according to their prevalence in the test sets, with metrics averaged over 100 runs to account for sampling variability. Since most cell subtypes are not identifiable from H&E alone by a pathologist, the task is challenging, and the random baseline provides a reference to show that our model outperforms chance-level expectations. In constrast, the upper bound baseline estimates the performance achievable with perfectly matched predictions at the cell level.
All models are trained on the OrionCRC training and validation splits to ensure a fair comparison. An overview of the compared models is shown in Figure 6. Pix2Pix, HEMIT, MIPHEI, and ROSIE are trained using their public implementations. As the DiffusionFT model of Oh et al. is not publicly available, we adapt the open Diffusion-E2E-FT implementation and use the same hyperparameters.
Evaluation is conducted on the OrionCRC test set (in-domain), and on HEMIT, PathoCell, Lizard, and PanNuke as out-of-domain datasets. Results across models and datasets are summarized in Tables 3 and 4, while Figure 7 presents the complete radar-plot comparison across datasets.
Figure 7: H&E-to-mIF Model Benchmark. Comparison of H&E-to-mIF models on OrionCRC (in-domain) and PathoCell, HEMIT, PanNuke, and Lizard (out-of-domain). We report cell-level AUPRC and F1 scores (cf Section 5.2), and pixel-level Pearson correlation. All models are trained on OrionCRC. As PanNuke and Lizard do not provide mIF data, only cell-level F1 scores are reported for these datasets
6.2.1 Benchmark Results
All models outperform prevalence-based random baselines, confirming that molecular information is to some degree predictable from morphology, as captured by H&E Images. Overall, MIPHEI-ViT consistently achieves the best performance across datasets and for nearly all markers. MIPHEI-ConvNeXt, ROSIE, and DiffusionFT follow with comparable results, with their relative ranking varying across datasets. DiffusionFT shows some performance drops on immune markers such as CD45 and CD45RO but performs well on expert-annotated datasets. HEMIT achieves strong in-domain performance but a substantial performance drop out of domain. Finally, Pix2Pix and nuclear morphometry baseline consistently perform worst.
Performance generally decreases when moving from in-domain to out-of-domain datasets. MIPHEI preserves a consistent best performane across datasets, highlighting its robustness. It is likely supported by the integration of the ViT foundation encoder. MIPHEI-ConvNeXt, HEMIT, and DiffusionFT have broadly similar behavior under domain shift. GAN-based approaches, including Pix2Pix and HEMIT, are strongly affected by domain shift, often producing near-zero predictions for non-nuclear markers. Overall, a large performance drop is observed on PathoCell for all models, including the upper bound. This is expected, as cell-type labels are derived from the full 58-marker panel while evaluation is restricted to the OrionCRC marker subset, which is not sufficient to recover all cell types.
In our ablation studies, we showed that MIPHEI-HEMIT, a variant using the HEMIT generator trained within the MIPHEI pipeline (no GAN, weighted MSE), achieves an overall cell-level AUPRC of 0.49 on the OrionCRC test set. Compared to the original HEMIT trained on the same data, this represents a substantial improvement of +0.10 AUPRC on the same split. This result demonstrates the effectiveness of our training pipeline, even when applied to an alternative architecture.
Model OrionCRC PathoCell HEMIT Lizard PanNuke
Nuclear Morphometry 0.247 0.111 0.357 0.352 0.393
Pix2Pix 0.287 0.097 0.311 0.235 0.305
HEMIT 0.390 0.094 0.375 0.227 0.345
ROSIE 0.433 0.187 0.644 0.408 0.397
DiffusionFT 0.392 0.169 0.588 0.451 0.498
MIPHEI-ConvNeXt (Ours) 0.437 0.197 0.577 0.467 0.462
MIPHEI-ViT (Ours) 0.517 0.250 0.647 0.517 0.60
Upper Bound 0.970 0.396 0.998
Table 3: Cell-level performance (AUPRC) across datasets and methods. Reported values correspond to AUPRC averaged across all markers and cell-level evaluation follows the protocol described in Section 5.2.2. MIPHEI achieves the highest AUPRC across all datasets. Dashes indicate metrics that are not applicable.
Model OrionCRC PathoCell HEMIT
Pix2Pix 0.153 0.038 0.141
HEMIT 0.280 0.028 0.245
ROSIE 0.354 0.222 0.479
DiffusionFT 0.324 0.228 0.513
MIPHEI-ConvNeXt (Ours) 0.404 0.278 0.517
MIPHEI-ViT (Ours) 0.470 0.336 0.538
Table 4: Pixel-level performance (Pearson correlation) across datasets and methods. Pearson correlation is computed over the full evaluation set (not averaged per tile) and averaged across all markers. MIPHEI achieves the highest Pearson correlation across datasets. Dashes indicate metrics that are not applicable.
6.2.2 Efficiency analysis.
Virtual staining typically requires running models over entire WSIs, making inference efficiency a critical factor. To inference compare computational cost, we report parameter counts, FLOPs, and peak GPU memory usage for a single 256×256 tile (Table 5). ROSIE is the most computationally demanding due to its sliding-window inference strategy. DiffusionFT is also costly, as the diffusion model is applied once per marker, even if a single inference timestep is used. MIPHEI-ViT is heavier than CNN-based models due to its ViT-H backbone, but thanks to LoRA, training involves only a small number of parameters. Pix2Pix and HEMIT are the most lightweight models but achieve lower accuracy. Overall, MIPHEI-ConvNeXt offers the best trade-off between efficiency and performance.
Model Params (M) FLOPs (G) VRAM (GiB)
Pix2Pix 54.4 6.49 0.23
HEMIT 79.7 65.17 0.40
ROSIE 49.5 8910.23 0.21
MIPHEI-ViT 1141.6 383.68 4.47
MIPHEI-ConvNeXt 203.3 49.49 0.82
DiffusionFT 951.3 6490.07 4.19
Table 5: Model inference computational summary. All results are reported for a $256\times 256$ input image using float32. Parameters are shown in millions (M), FLOPs in gigaFLOPs (G), and VRAM in GiB. Only 6.7M parameters are trainable.
@@ 6.3 Marker-Level Analysis
In this section, we focus on the analysis of MIPHEI. Results are illustrated in Figure 8.
Figure 8: MIPHEI Evaluation analysis: (a) Marker associations and phenotypic groupings used to interpret differences in marker predictability. Arrows denote dominant subset relationships between phenotypes (e.g., CD3e+ cells largely co-express CD45). (b) Cell-level performance of MIPHEI on OrionCRC. Mean predicted marker expression is extracted per cell and used to train a logistic regression classifier to match pseudo cell types on validation cells, then evaluated on test cells. We report cell-level AUPRC and F1 scores, with a prevalence-based random baseline for F1 to assess whether meaningful biological patterns are learned. Error bars indicate 95% confidence intervals obtained by bootstrapping (Section 5.2). (c) IMMUcan evaluation of MIPHEI on consecutive sections. Pearson correlation and linear regression plots between predicted and pseudo-labeled cell-type counts across 17k tiles.
@@ 6.3.1 In-domain performance analysis
Markers with the highest performance include epithelial markers E-cadherin (F1 0.92) and Pan-CK (F1 0.93), which label epithelial cells forming well-defined clusters and glandular structures in H&E. The immune marker CD45 (F1 0.70) also performs well. While it mainly identifies lymphocytes, which are easily recognizable in H&E, it also includes cells like monocytes, which are harder to spot, making prediction a bit more challenging. $\alpha$-SMA also achieves strong performance (F1 0.77), as it labels stromal and vascular structures that can be identifiable on H&E.
Markers with moderate performance included CD31 (F1 0.37), which mark endothelial cells forming vascular structures visible on H&E. However, these cells represent a relatively small population at the cell level, which makes accurate cell-level prediction more challenging. Immune subtype markers CD3e, CD45RO, CD4 and CD20 (F1 0.29–0.63) also show moderate performance, with broader T-cell markers like CD3e (F1 0.63) achieving better performance than more specific ones like CD45RO (F1 0.56). CD20, a B-cell marker, shows lower performance within this group, likely due to the more limited representation of B cells in OrionCRC. While lymphocytes are visible in H&E, their subtypes remain indistinguishable to pathologists, highlighting our model’s value on this difficult task. Macrophage markers CD68 (F1 0.362) and CD163 (F1 0.206) faced similar challenges, as their heterogeneity complicates identification, with CD163 being a macrophage subtype of CD68, further complicating prediction. Ki-67 also exhibits moderate performance, as it labels proliferating cells across the G1, S, and G2 phases that lack distinctive morphological features on H&E, with only a small fraction corresponding to visible mitotic figures. Finally, CD68 shows moderate performance, likely due to the heterogeneous morphology of macrophages and the diffuse nature of its expression.
Markers with the lowest performance are CD163 (F1 0.27), FOXP3 (F1 0.22), CD8a (F1 0.19), and PD-L1 (F1 0.05). CD163 shows limited performance because it marks M2 macrophages, a functional subset of CD68+ cells, and H&E probably does not allow to distinguish macrophage activation states. FOXP3 labels rare regulatory T-cells, a highly specific CD4+ subset, making it one of the most difficult immune markers to predict from morphology alone. Similarly, CD8a marks a less abundant T-cell subtype that is typically more dispersed than CD4+ cells. Finally, PD-L1 is a functional marker that is rarely expressed and not associated with a specific cell type, and shows low-intensity, noisy staining in OrionCRC, making prediction from H&E particularly challenging.
In summary, epithelial markers, broad immune markers and $\alpha$-SMA seem predictable with high accuracy. Markers defining specific immune subtypes are harder to predict, and some functional markers are not predictable with an accuracy high enough to be applicable in practice. Overall, prediction performance decreases as markers correspond to increasingly rare or specialized cell populations.
@@ 6.3.2 Out-of-domain performance analysis
Our external validation on HEMIT, PathoCell and IMMUcan confirms that MIPHEI, trained on OrionCRC, generalizes well to datasets with domain shifts in mIF technology and H&E appearance, requiring minimal adaptation.
On the HEMIT dataset, analysis is limited by the small number of available markers (CD3e and Pan-CK), yet MIPHEI shows good performance with minimal adaptation. These results should be interpreted with caution, as tiles were selected from regions containing many CD3+ lymphocytes and Pan-CK+ cells, which simplifies the task and may inflate performance. In this setting, a model predicting all lymphocytes as T cells would already perform well. At the same time, stain normalization applied to all H&E images introduces a strong, artificial domain shift.
On PathoCell, overall performance is moderate, including for the upper bound, due to the mismatch between the full 58-marker PathoCell panel used for labeling and the restricted set of predicted markers, as discussed earlier. Despite this limitation, we observe trends consistent with the in-domain analysis, with the best performance for T cells, B cells, smooth muscle, and tumor cells, which are directly linked to the predicted protein markers. Although absolute scores decrease, the gap relative to the upper bound remains similar. MIPHEI even outperforms the upper bound for granulocytes, likely because the predicted panel lacks a granulocyte-specific marker and the models leverage morphological proxies. At the pixel level, Hoechst reconstruction remains strong but other marker degrades compared to OrionCRC. The absence of autofluorescence subtraction, which slightly shifts the target signal, may explain this behavior.
We further evaluate model generalization on the IMMUcan dataset by correlating cell-type counts predicted by MIPHEI from H&E with pseudo-label counts obtained from consecutive mIF sections. Pearson correlation is computed over 0.26 mm2 regions using a logistic regression cell classifier trained on OrionCRC validation cells with MIPHEI single-cell predictions. We observe strong correlations for CD4 (0.80), CD3e (0.73), and CD8 (0.69), and moderate correlations for FOXP3 (0.60) and Pan-CK (0.63). We also observe overdetection for CD3e (i.e. cells with CD3e prediction and measured 0 expression) and underdetection for Pan-CK (cells with predicted 0-expression and positive according to the measurement). Manual inspection showed that this was mainly due to low-quality tiles with high auto-fluorescence. In Figure 9.A, we observe a similar case on another dataset, with an artifact present in the target but correctly not predicted by the model.
On Lizard, performance is good for epithelial cells, and remains acceptable for connective tissue ($\alpha$-SMA) and lymphocytes. Plasma cells and neutrophils are weakly supported by the predicted marker panel and therefore show low performance. Eosinophils, which are not specifically covered by the panel, exhibit very low performance. On PanNuke, the cross-organ setting and higher magnification introduce a stronger domain shift, leading to overall lower performance. Neoplastic and dead cell classes show more limited performance, as their identification relies on indirect cues from the predicted panel. Neoplastic cells can be inferred from Pan-CK, E-cadherin, and Ki67, while dead cells are characterized by low or absent signal across all predicted markers. Unexpectedly, connective tissue performance is also limited, which may be partly explained by the very small number of positive cells.
Figure 9: WSI prediction of mIF from H&E using MIPHEI-ViT. Top left: MIPHEI-ViT prediction overlay (RGB composite of 16 mIF markers) for a full CRC WSI. Bottom left: corresponding real mIF target. Right: zoom-in on four selected regions (A–D) with model predictions (top) and ground truth (bottom). The regions cover diverse tissue contexts, including an artifact in (A) that is correctly ignored by the model. (see Supplementary Video 2 for an interactive side-by-side demonstration in QuPath).